OpenStudy (anonymous):

Which expression is the simplified form of sqrt(72x^5y^4)

OpenStudy (cwrw238):

sqrt 72 = sqrt (36 * 2) = sqrt36*sqrt2 = 6 sqrt2 can you understand that?

OpenStudy (anonymous):

i understand that part

OpenStudy (cwrw238):

and the square roots of x^5 and y^4 are found by dividing the indices by 2

OpenStudy (cwrw238):

eg sqrt x^5 = x^(5 /2)

OpenStudy (anonymous):

you lost me

OpenStudy (cwrw238):

y^4 = y^2 * y^2 right? (by the law of indices)

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

y^a * y^b = y^(a+b)

OpenStudy (cwrw238):

so y^2 is square root of y^4

OpenStudy (cwrw238):

in similar way y^(5/2) * y(5/2) = y^5 to get the square root we divide the index ( or power) by 2